Wavelength of a wave is 1,500m calculate its corresponding frequency.
Answer #1

Given that,

\lambda = 1500 m

We know that,

Frequency f = v / \lambda

take speed of wave, v = 3*10^8 m/s

f = 3*10^8 / 1500

f = 2*10^5 Hz